OpenStudy (anonymous):

what are two consecutive numbers add to 257

OpenStudy (anonymous):

774

OpenStudy (anonymous):

got that?

OpenStudy (anonymous):

257+258+259=774

OpenStudy (anonymous):

two consecutive numbers that add up to 257 means you have x + (x+1) = 257 solve for x

OpenStudy (judygreeneyes):

Exactly, so where did 774 come from? So you want 128 and 129

OpenStudy (anonymous):

the trick is simple first substract by 1 and divide by 2 then you got the first answer then add it with 1 then you got the second answer Tada!

OpenStudy (anonymous):

(257-1)/2 = 128 add with 1 = 129

OpenStudy (kainui):

You have a number, let's call it N. Now what's the number after N? That's our consecutive number, which is just whatever N is, plus 1. So here we have: N and (N+1) and supposedly if you add these two numbers together you get 257. So let's just do that: N+(N+1)=257 Now you can just do a little algebra to solve for N. I hope that explanation makes it more clear, even though it seems weird. Try plugging in a number for N before solving to see why N+1 is a consecutive number. So for example, if you have N and N+1 then if you plug in 7 for N, you will see that you get 7 and 8.

OpenStudy (anonymous):

thx guys helped a lot!

OpenStudy (anonymous):

its like my 2week of year7 and wasn't sure how to do it!

OpenStudy (campbell_st):

just go 257/2 = 128.5 so the 2 numbers are the integers either side of 128.5 so the numbers are 128 and 129

OpenStudy (kainui):

I like that answer @campbell_st since it's clever, but unfortunately it's not as helpful if you wanted to find out what numbers you have when you have the sum of say, 5 consecutive numbers.